The width of a rectangle is 4 inches shorter than its length. The perimeter of the rectangle is 488 inches. What are the length and width of the rectangle?
Found 2 solutions by swincher4391, shweta:
Answer by swincher4391(1107)

By definition, perimeter = 2*length + 2*width.
Let the length = l, then the width = l-4
So 2*l + 2(l-4) = 488
2l + 2l - 8 = 488
4l = 456
l = 114
w = 110

Answer by shweta(56)

Let the width of a rectangle be 'w' and length be 'l'
Given: w=l-4 .....(1)
and perimeter,P=488 inches
Formula for perimeter of a rectangle= 2(l+w)
2(l+w)= 488
Now put in the value of 'w' in the above formula
2(l+l-4)=488
2(2l-4)=488
Now open the bracket
2*2l-2*4=488
4l-8=488
4l=488+8
4l=496
l=496/4
l=124
Length of the rectangle is 124 inches
Substitute the value of 'l' in equation 1
w=l-4
w=124-4
w=120
The width of the rectangle is 120 inches